Circular No. 5634

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1992 No. 3 Paul J. Camilleri, Cobram, Victoria, Australia, reports his photo-

graphic discovery (using an 85-mm camera lens + T-Max 400 film) of an apparent nova, providing the following photovisual magnitude estimates: Sept. 29.41 UT, [12; Oct. 13.40, 9.0. Camilleri notes that there is no image in the immediate position of the nova on Union Observatory prints from 1927 Aug. 19, to a limiting magnitude of 15.

R. H. McNaught, Anglo-Australian Observatory, confirms the apparent nova from a pair of CCD images taken with the 1.0-m reflector at Siding Spring on Oct. 13.54. The nova’s position is α = 18^{h}20^{m}37^{s}, δ = −23^{o}01′.4 (equinox 1950.0). The object lies ∼ 1′ east and 3′ north of SAO 186756. The brightest possible candidate on the ESO B and R and the SERC J surveys is a star of mag ∼ 18.

SUPERNOVA 1992bb IN IRAS 21156–0747 M. Della Valle and R. Corradi, European Southern Observatory, re-

port: “A spectrogram (range 490-690 nm, resolution ∼ 1 nm) of SN 1992bb (cf. IAUC 5629) was obtained with the 1.5-m telescope (+ Boller & Chivens spectrograph) at La Silla on Oct. 6.2 UT. Preliminary analysis reveals this object to be a type-Ia supernova. The expansion velocity deduced from the minimum of the Si II 635-nm absorption is ∼ 10 000 km/s.”

PERIODIC COMET SWIFT-TUTTLE (1992t) Total visual magnitude estimates (cf. IAUC 5627): Oct. 2.05 UT, 9.0

(G. Kronk, Troy, IL, 0.33-m reflector); 3.50, 8.6 (C. S. Morris, Pine Mountain Club, CA, 20×80 binoculars); 4.37, 8.4 (D. W. E. Green, near Franklin, NH, 12×50 binoculars); 6.39, 8.6 (J. E. Bortle, Stormville, NY, 20×80 binoculars); 7.39, 8.2 (Bortle); 7.50, 8.3 (C. S. Morris, near Mt. Wilson, CA, 20×80 binoculars); 8.15, 8.0 (R. Haver, Viterbo, Italy, 15×80 binoculars); 9.51, 8.0 (Morris); 12.98, 7.9 (Bortle).

NOVA CYGNI 1992 Photometry by R. Monella, Covo, Italy: Sept. 27.892 UT, V = 9.29,

B–V = +0.17; 30.934, 9.32, +0.16; Oct. 2.967, 9.32, +0.17 (comparison star BD +52^{o}2734).

1992 October 13 Daniel W. E. Green